Find the x- and y- intercepts of the line y=-2x-2.
Answer by stanbon(75887)

Find the x- and y- intercepts of the line y=-2x-2
--------
x-intercept ?
Let y = 0
Solve for "x"::
0 = -2x-2
x = 2/-2 = -1
--
x-int is (-1,0)
-----------
y-intercept = f(0) = -2*0-2 = -2
-----------
y-int is (0,-2)
-----------------
Cheers,
Stan H.
------------